suppose that the number of bacteria in a certain population increases according to a continuous exponential growth model. a sample of 2500 bacteria selected from the population reached the size of 2770 bacteria in 4 hours. find the hourly growth rate parameter
Answer by stanbon(75887)

suppose that the number of bacteria in a certain population increases according to a continuous exponential growth model. a sample of 2500 bacteria selected from the population reached the size of 2770 bacteria in 4 hours. find the hourly growth rate parameter
-----------------
y = a*b^t
------
2770 = 2500*b^4
----
b^4 = 1.108
---
b = 1.2597
-----------------
y = 2500*(1.2597)^t
-----
Growth rate: # of bacteria increases at the rate of 26% per hour
-----
Cheers,
Stan H.
--------------